### v1

@@ 4.2 MFGDiffusion
Figure 3: The network architecture of MFGDiffusion. Additional mask and masked image features interact with the intermediate features of U-Net through cross-attention and are subsequently injected into the next layer of U-Net.
4.3 Fusion Process
@@
Although both the feature extractor and the U-Net architecture are based on CNNs, the feature dimensions are $(B,C_{1},H_{1},W_{1})$ , while the intermediate representations of the U-Net have dimensions $(B,C_{2},H_{2},W_{2})$ Therefore, to inject the features, we need to project the channel dimensions of the mask and masked image to match the dimensions of the U-Net’s intermediate representations using a Linear layer. After projection, we used cross-attention to fuse the features of the mask and masked image with the corresponding U-Net layers and input them into the next layer. For the intermediate layers where both the mask and masked image features need to be injected simultaneously, we designed a joint cross-attention mechanism to fuse the features of both with the output of the corresponding U-Net layer, as shown in Figure 4. We first perform an attention fusion of $X$ with the mask feature $F_{m}$ and the masked image feature $F_{M}$ separately,resulting in two intermediate representations. These representations are then concatenated and passed through an MLP to enhance their fusion. In the final output, we also employed residual connections [41] to improve the effectiveness of the attention mechanism. The final output serves as the input to the next layer of the U-Net. Formally, this joint cross-attention process can be expressed as follows:
@@
Figure 4: Joint Cross-Attention for features. Joint Cross-Attention for features: Simultaneously inject the features of both the mask and the masked image through joint cross-attention to collaboratively guide the denoising process.
4.4 Mask Random Difference Loss
@@
We continued to employ Mean Squared Error (MSE) to compute the loss, where $\omega$ represents the weight of the Mask Random Difference Loss, and $C_{all}$ denotes the conditions guiding the denoising process, including text, mask, and masked image. In Section 5.3, we designed an ablation study for the Mask Random Difference Loss, ultimately setting $\omega=0.4$ as the parameter for model training.
4.5 Synthetic Data Filtering
During the data synthesis phase, since our ultimate goal is to generate plausible smoke in smoke-free background images while leveraging the text-guided capabilities of the pre-trained model, we require captions for smoke-free background images that include descriptions of smoke to guide the smoke generation process. Here, we utilized a large language model to directly generate smoke-inclusive captions from the captions of smoke-free images. Owing to the robust capabilities of the large language model, it typically incorporates smoke-related information into the captions of smoke-free background images in a reasonable manner, such as modifying sentence structures by adding phrases like "with smoke," "a smoke plume in," or "add smoke to." Given that the scenes are forest environments with relatively simple background information, we do not need to exert excessive control over the textual details; instead, we relied on the fundamental capabilities of the text within the pre-trained model. For the generation of synthetic smoke, we used the segmentation results from SAM as the mask input, with specific details discussed in Section 5.1.
@@ 4.5 Synthetic Data Filtering
Due to the widespread application of diffusion model-based image generation techniques, many datasets for inpainting models (Brooks et al. (2023), Qi et al. (2024), Wei et al. (2024)) utilize existing image generation technologies to produce data for their specific tasks. In previous tasks involving the generation of datasets, most researchers have employed ClipScore (Radford et al. (2021)) to evaluate text-image consistency, thereby filtering the generated data. However, prior work (Jiang et al. (2024)) has shown that ClipScore’s assessment of text-image consistency has a low correlation with human judgment. Therefore, drawing inspiration from recent work (Wei et al. (2024)), we utilized a multimodal large language model to evaluate the generated data. Unlike previous approaches, since the smoke data is generated from smoke-free backgrounds primarily featuring forest scenes, evaluating text-image consistency would be influenced by the similarity of the backgrounds. Thus, we utilized the multimodal large language model to assess the state of the generated smoke rather than text-image consistency for data filtering. Figure 5 illustrates the overall workflow of our data filtering process.
Although we can directly employ a multimodal large language model to assess the quality of smoke generation, this approach has several limitations. For example, the multimodal large language model is prone to misjudgment when directly evaluating smoke, and the generated evaluation scores tend to be closely clustered, lacking sufficient granularity. Such a filtering method is ineffective in significantly enhancing the quality of the synthetic dataset. Therefore, we first need to establish an evaluation framework for smoke. To simplify our evaluation system, we focused solely on potential defects in the smoke generation process and the inherent characteristics of smoke. Ultimately, we selected three evaluation metrics: smoke color, visibility, and semi-transparency. Each metric is scored on a scale of 0 to 10. Although the smoke metrics have been specified, testing revealed that the multimodal large language model still exhibits clustering tendencies in score evaluation, making it difficult to distinguish between superior and average generation results. To obtain a more effective data filter, we manually annotated 100 smoke data points, including scores for the three aforementioned evaluation metrics. We then fine-tuned Qwen2-VL using this annotated data and employed the fine-tuned Qwen2-VL to filter our dataset. Finally, we selected the top 50% of the data as our final synthetic dataset, which was combined with the real dataset and used to train the detection model.
@@
5 EXPERIMENTS
5.1 Implementation Details
Training Details.We employed Stable Diffusion 2 inpainting as our foundational inpainting model, which consists of a total of 4 downsampling modules, 4 upsampling modules, and an intermediate linking layer. These layers are numbered from 0 to 8, where layers 0-3 represent the downsampling process from high-resolution to low-resolution, layer 4 is the intermediate layer, and layers 5-8 represent the upsampling process from low-resolution to high-resolution. Based on the characteristics of each layer, we ultimately chose to inject the image features of the masked image into layers 0 and 8, while simultaneously injecting the features of both the mask and masked image into layers 1 and 7. The intermediate layer 4 is exclusively injected with mask features. The cross-attention modules used in each layer that receives feature injection do not share weights and independently guide the denoising process. During training, we froze the parameters of the VAE, U-Net, text encoder, and the pre-trained feature extraction module (ResNet50), updating only the projection layers after feature output and the cross-attention modules used for injecting features into the corresponding layers of the denoising U-Net. The weight for the Mask Random Difference Loss is set to 0.4. The models are trained on a single A100-80G GPU with a total batch size of 32. We adopted AdamW (Loshchilov (2017)) as the optimizer with a learning rate of $1\times 10^{-4}$ and trained for a total of 20,000 iterations. For inference, we used classifier-free guidance with a guidance scale of 7.5 and perform 50 denoising steps.
@@ 5.1 Implementation Details
Data filtering details. We utilized the pre-trained multimodal large language model, Qwen2-VL-2B-Instruct, as our data filtering model. As described earlier, we first manually annotated 100 smoke data points, assigning scores for three specific characteristics. However, manual scoring inherently involves subjective factors, so we established certain scoring principles to ensure more objective annotations. For instance, fully transparent smoke is constrained to a score of 8-10 for semi-transparency; poorly generated smoke colors that blend with the background or appear overly white are constrained to around 5; and images where smoke generation is not clearly visible are constrained to scores of 0-2 across all three metrics. Nevertheless, due to the diversity of generated smoke data, manual scoring can generally achieve a relatively objective assessment in most cases. Through this process, we obtained evaluation data for 100 images. We then used these 100 data points to fine-tune the advanced multimodal large language model, Qwen2-VL. After fine-tuning, Qwen2-VL gained the capability to evaluate smoke color, visibility, and semi-transparency. Based on empirical knowledge of the quality of generated smoke data, we weighted the scores of the three characteristics to calculate a final score, assigning weights of 0.5, 0.3, and 0.2 to color, visibility, and semi-transparency, respectively. Using Qwen2-VL, we scored the generated smoke images, performed weighted ranking, and selected the top 50% of the synthetic smoke data as the final synthetic dataset. A partial display of our synthetic dataset is shown in Figure 7.
5.2 Results
@@
Qualitative evaluation. We compared the most advanced inpainting models that utilize masks and text as conditions, including Stable Diffusion V2.1 (Rombach et al. (2022)), which modifies strategies based on local masks, Blended Latent Diffusion (BLD) (Avrahami et al. (2023)), the end-to-end inpainting model Stable Diffusion Inpainting (SD-Inpainting), and the versatile editing model PowerPaint (Zhuang et al. (2025)). To ensure fairness in qualitative comparison, we fine-tuned the base models of these methods and then applied their respective strategies or operations to generate smoke images. As shown in Figure 8, our method demonstrates superior performance compared to previous approaches. The generated smoke exhibits more reasonable results in terms of color, visibility, and semi-transparency, without issues such as excessive smoke density, distortion, or generation failure. Moreover, our method showcases strong capabilities in preserving the semi-transparent properties of smoke edges and maintaining background integrity, enabling seamless interaction with the background. In terms of generation quality, our method can produce images with a resolution of $1024\times 1024$, providing the smoke detection model with richer details of both the target and the background.
@@
Table 1: Comparison of generative models. Our method achieves state-of-the-art or competitive results across various metrics for smoke image generation.
@@
(1) Metrics for image inpainting tasks. Unlike some previous works, we did not use FID (Fréchet Inception Distance) as an evaluation metric for image quality, as it is unable to accurately assess the diverse content generated by current diffusion models. Instead, we adopted commonly used metrics for image generation tasks as our evaluation criteria, including: Peak Signal-to-Noise Ratio (PSNR) (Sheikh et al. (2006)), Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al. (2018b)), Mean Squared Error (MSE), and Structural SIMilarity (SSIM) (Wang et al. (2004)). Additionally, we employed CLIP Similarity (CLIP Sim) (Radford et al. (2021)) to evaluate the similarity between the generated smoke and the text, thereby assessing the realism of the smoke and the quality of the generated images.
(2) Metrics for object detection tasks. We used four metrics to evaluate the performance of models trained on both real and mixed datasets:: Precision (Prec), Recall, mAP50, and mAP50-95. Precision measures the proportion of correctly predicted positive instances among all predicted positive instances. Recall measures the proportion of actual positive instances that are correctly identified. mAP50 is used to evaluate the accuracy of object localization by the model at different confidence levels, considering only predictions with an Intersection over Union (IoU) greater than or equal to 0.5. mAP50-95 is an extended version of mAP50, which not only calculates the Average Precision (AP) at an IoU threshold of 0.5 but also computes the AP at multiple IoU thresholds ranging from 0.5 to 0.95 in increments of 0.05, ultimately averaging these AP values to obtain mAP50-95. These metrics help us assess the improvement in the detection model’s performance due to the synthetic data.
Table 2: Comparison of different detection models. By combining real data with synthetic data, we significantly improved the performance of the smoke detection model.
@@
Table 1 presents the comparison results between our method and previous works, demonstrating the effectiveness of MGFDiffusion in terms of image quality and text alignment. The modified Stable Diffusion, which employs a strategy-based approach, exhibits the poorest image quality. This is because the smoke in the images is generated entirely based on the mask, without considering the interaction between the smoke edges and the background, resulting in either failed smoke generation or overly dense and distorted smoke. PowerPaint performs poorly in generating semi-transparent objects, which we attribute to its training being overly biased toward generating objects with fixed shapes. Simply fine-tuning SD-Inpainting is insufficient to fully utilize the information from the mask and masked image, making it difficult to generate reasonably visible smoke within the mask and causing interference from the model’s prior knowledge. Blended Latent Diffusion shows unstable generation performance, failing to produce smoke in certain scenarios.
Table 3: Comparison of synthetic data and real data. With an equivalent amount of data, the inclusion of our synthetic data results in more substantial performance gains for the smoke detection model compared to the addition of real data.
@@
Yolov6 0.637 0.33 0.641 0.642 0.759 0.464 0.781 0.677
Yolov8 0.675 0.35 0.71 0.665 0.795 0.495 0.79 0.728
@@
Yolov10 0.67 0.341 0.737 0.636 0.777 0.483 0.815 0.688
Yolov11 0.68 0.349 0.647 0.725 0.793 0.495 0.811 0.721
In Table 2, we demonstrated the effectiveness of our synthetic dataset. We conducted experiments using widely adopted detection models, including Yolov6, Yolov8, Yolov9, Yolov10, and Yolov11. To evaluate the performance of the trained detection models in real-world smoke detection scenarios, we used only real data as the test set. The experimental results indicate that mixing the synthetic dataset with the real dataset leads to significant improvements across all metrics for each detection model. This confirms that our high-quality synthetic dataset can effectively enhance the performance of smoke detection models in real-world environments.
We also compared the effects of replacing the synthetic data with real data during model training, as shown in Table 3. It can be observed that, with the same amount of data, incorporating synthetic data leads to greater improvements in the smoke detection model compared to adding real data. We attributed this to the fact that the synthetic smoke data is inherently very close to real data, and through the data filtering process, only high-quality synthetic data is retained. Therefore, adding this portion of high-quality synthetic data to the real dataset yields better results than simply adding more real data.
Figure 9: Ablation of modules in network architecture. Using only the mask can lead to a decrease in the ability to interact with the background, while using only the masked image or not employing the Joint Cross-Attention (JCA) mechanism can result in varying degrees of degradation in the generation capability.
5.3 Ablation Study
Figure 10: Ablation of feature extractor. ResNet50 serves as a more stable feature extractor and yields better generation outcomes.
Figure 11: Ablation of mask random difference loss. Varying the weights can significantly influence the model’s capability to effectively interact with the background.
Table 4: Ablation of filtering module. The filtering module significantly enhances data quality, and utilizing this high-quality data for training leads to a smoke detection model with superior performance.
@@ 5.3 Ablation Study
We also designed experiments to validate the effectiveness of the multimodal large language model in data filtering. For this purpose, we created two datasets: one using filtered synthetic smoke data and the other using randomly selected synthetic smoke data, both mixed with real data in a 1:1 ratio, with a positive-to-negative sample ratio of 1:1. We trained multiple detection models on these two datasets separately. As shown in Table 4, significant improvements were observed across all four metrics used in our evaluation for different detection models.
6 Conclusion
In this paper, we proposed a comprehensive pipeline for the generation and detection of forest fire smoke images. This includes the acquisition of a pre-trained diffusion model dataset, a feature extraction module, a novel auxiliary loss function, and a data filtering module based on a multimodal large language model. We utilized various pre-trained multimodal models to generate a dataset of over 60,000 samples for training the diffusion model, and we trained our designed projection layer and cross-attention module in conjunction with the Mask Random Difference Loss. Using the proposed new model architecture, we generated smoke data for smoke-free backgrounds. We found that the designed new model architecture was capable of generating sufficiently realistic forest fire smoke images, surpassing previous methods in terms of visual effects such as smoke texture, morphology, and semi-transparency. Additionally, it exhibits more reasonable consistency with the background. The generated smoke images outperform existing models on metrics such as PSNR, SSIM, LPIPS, MSE, and ClipSim.
Furthermore, based on the characteristics of the generated smoke, we fine-tuned and utilized a multimodal large language model to assist in filtering high-quality smoke datasets. Combining real data with synthetic data effectively addresses challenges such as the difficulty in acquiring smoke datasets and their limited diversity, while also enhancing the performance of smoke detection models. Empirical evidence demonstrates that augmenting smoke datasets with synthetic data significantly improves model capabilities, making it a worthwhile area for further investigation. Our future research will focus on generating fire smoke in a broader range of scenarios and creating video content of smoke scenes. Given the rapid advancements in video generation technology, we will consider the dynamic motion of smoke as a key aspect of our subsequent work.
Acknowledgements
This work was financially supported by the National Natural Science Foundation of China under Grant No. 32471866, the Anhui Provincial Science and Technology Major Project under Grant No. 202203a07020017, the Central Government-Guided Local Science and Technology Development Funds of Anhui Province under Grant No. 2023CSJGG1100 and the Key Project of Emergency Management Department of China under Grant No. 2024EMST010101. This research was also supported by the advanced computing resources provided by the Supercomputing Center of the USTC.The authors gratefully acknowledge all of these supports.

### v2

@@ IV-B MFGDiffusion
Figure 4: Joint Cross-Attention for features. Joint Cross-Attention for features. Simultaneously inject the features of both the mask and the masked image through joint cross-attention to collaboratively guide the denoising process. The two attention outputs are fused and combined with residual connections to update the U-Net representation.
IV-C Fusion Process
@@
Although both the feature extractor and the U-Net architecture are based on CNNs, the feature dimensions are $(B,C_{1},H_{1},W_{1})$ where $B$ is the batch size, $C_{1}$ is the number of channels, $H_{1}$ is the feature map height, and $W_{1}$ is the feature map width,while the intermediate representations of the U-Net have dimensions $(B,C_{2},H_{2},W_{2})$ with $C_{2}$ , $H_{2}$ , $W_{2}$ defined similarly. Therefore, to inject the features, we need to project the channel dimensions of the mask and masked image to match the dimensions of the U-Net’s intermediate representations using a Linear layer. After projection, we used cross-attention to fuse the features of the mask and masked image with the corresponding U-Net layers and input them into the next layer. For the intermediate layers where both the mask and masked image features need to be injected simultaneously, we designed a joint cross-attention mechanism to fuse the features of both with the output of the corresponding U-Net layer, as shown in Figure 4. We first perform an attention fusion of $X$ with the mask feature $F_{m}$ and the masked image feature $F_{M}$ separately,resulting in two intermediate representations. These representations are then concatenated and passed through an MLP to enhance their fusion. In the final output, we also employed residual connections [41] to improve the effectiveness of the attention mechanism. The final output serves as the input to the next layer of the U-Net. Formally, this joint cross-attention process can be expressed as follows:
@@
IV-D Mask Random Difference Loss
@@
We continued to employ Mean Squared Error (MSE) to compute the loss, where $\omega$ represents the weight of the Mask Random Difference Loss, and $C_{all}$ denotes the conditions guiding the denoising process, including text, mask, and masked image. In Section V-C, we designed an ablation study for the Mask Random Difference Loss, ultimately setting $\omega=0.4$ as the parameter for model training.
IV-E Synthetic Data Filtering
During the data synthesis phase, since our ultimate goal is to generate plausible smoke in smoke-free background images while leveraging the text-guided capabilities of the pre-trained model, we require captions for smoke-free background images that include descriptions of smoke to guide the smoke generation process. Here, we utilized a large language model to directly generate smoke-inclusive captions from the captions of smoke-free images. Owing to the robust capabilities of the large language model, it typically incorporates smoke-related information into the captions of smoke-free background images in a reasonable manner, such as modifying sentence structures by adding phrases like ”with smoke,” ”a smoke plume in,” or ”add smoke to.” Given that the scenes are forest environments with relatively simple background information, we do not need to exert excessive control over the textual details; instead, we relied on the fundamental capabilities of the text within the pre-trained model. For the generation of synthetic smoke, we used the segmentation results from SAM as the mask input, with specific details discussed in Section V-A.
@@ IV-E Synthetic Data Filtering
Due to the widespread application of diffusion model-based image generation techniques, many datasets for inpainting models [32], [33], [34] utilize existing image generation technologies to produce data for their specific tasks. In previous tasks involving the generation of datasets, most researchers have employed ClipScore [26] to evaluate text-image consistency, thereby filtering the generated data. However, prior work [36] has shown that ClipScore’s assessment of text-image consistency has a low correlation with human judgment. Therefore, drawing inspiration from recent work [34], we utilized a multimodal large language model to evaluate the generated data. Unlike previous approaches, since the smoke data is generated from smoke-free backgrounds primarily featuring forest scenes, evaluating text-image consistency would be influenced by the similarity of the backgrounds. Thus, we utilized the multimodal large language model(MLLM) to assess the state of the generated smoke rather than text-image consistency for data filtering. Figure 6 illustrates the overall workflow of our data filtering process.
While MLLM can be directly used to evaluate the quality of generated smoke, this approach has notable limitations: the model is prone to misjudgment, and the scores tend to cluster with insufficient differentiation, making it ineffective for significantly improving the quality of the synthetic dataset. To address this, we first established a simplified evaluation framework focusing on defects in smoke generation and inherent smoke characteristics, including three metrics—color, visibility, and semi-transparency—each scored on a 0–10 scale. Testing revealed that MLLM scores still exhibited clustering tendencies. Therefore, we manually annotated 100 sets of smoke data to fine-tune Qwen2-VL and used the fine-tuned model to filter the data, ultimately selecting the top 50% of scored samples. These selected synthetic data were combined with real data to train the detection model.
@@
V EXPERIMENTS
V-A Implementation Details
Training Details.We employed Stable Diffusion 2 inpainting as our foundational inpainting model, which consists of a total of 4 downsampling modules, 4 upsampling modules, and an intermediate linking layer. These layers are numbered from 0 to 8, where layers 0-3 represent the downsampling process from high-resolution to low-resolution, layer 4 is the intermediate layer, and layers 5-8 represent the upsampling process from low-resolution to high-resolution. Based on the characteristics of each layer, we ultimately chose to inject the image features of the masked image into layers 0 and 8, while simultaneously injecting the features of both the mask and masked image into layers 1 and 7. The intermediate layer 4 is exclusively injected with mask features. The cross-attention modules used in each layer that receives feature injection do not share weights and independently guide the denoising process. During training, we froze the parameters of the VAE, U-Net, text encoder, and the pre-trained feature extraction module (ResNet50), updating only the projection layers after feature output and the cross-attention modules used for injecting features into the corresponding layers of the denoising U-Net. The weight for the Mask Random Difference Loss is set to 0.4. The models are trained on a single A100-80G GPU with a total batch size of 32. We adopted AdamW [37] as the optimizer with a learning rate of $1\times 10^{-4}$ and trained for a total of 20,000 iterations. For inference, we used classifier-free guidance with a guidance scale of 7.5 and perform 50 denoising steps.
@@ V-A Implementation Details
Data filtering details. We utilized the pre-trained multimodal large language model, Qwen2-VL-2B-Instruct, as our data filtering model. As described earlier, we first manually annotated 100 smoke data points, assigning scores for three specific characteristics. However, manual scoring inherently involves subjective factors, so we established certain scoring principles to ensure more objective annotations. For instance, fully transparent smoke is constrained to a score of 8-10 for semi-transparency; poorly generated smoke colors that blend with the background or appear overly white are constrained to around 5; and images where smoke generation is not clearly visible are constrained to scores of 0-2 across all three metrics. Nevertheless, due to the diversity of generated smoke data, manual scoring can generally achieve a relatively objective assessment in most cases. Through this process, we obtained evaluation data for 100 images. We then used these 100 data points to fine-tune the advanced multimodal large language model, Qwen2-VL. After fine-tuning, Qwen2-VL gained the capability to evaluate smoke color, visibility, and semi-transparency. Based on empirical knowledge of the quality of generated smoke data, we weighted the scores of three key characteristics—color, visibility, and semi-transparency—to calculate a final score, assigning weights of 0.5, 0.3, and 0.2, respectively. This weighting scheme is motivated by both domain knowledge and empirical observations. For smoke detection, color provides the most direct cue for distinguishing smoke from the background, and deviations in color severely degrade detection accuracy; therefore, it receives the highest weight. Visibility determines whether smoke can be reliably captured by the model, and insufficient visibility often leads to missed detections, making it the second most important factor. Semi-transparency, while a natural attribute of smoke, has a relatively smaller impact on model performance once color and visibility are within reasonable ranges, and is therefore assigned the lowest weight.Using Qwen2-VL, we scored the generated smoke images, performed weighted ranking, and selected the top 50% of the synthetic smoke data as the final synthetic dataset. A partial display of our synthetic dataset is shown in Figure 7.
V-B Results
@@
Qualitative evaluation. We compared the most advanced inpainting models that utilize masks and text as conditions, including Stable Diffusion V2.1 [23], which modifies strategies based on local masks, Blended Latent Diffusion (BLD) [38], the end-to-end inpainting model Stable Diffusion Inpainting (SD-Inpainting), and the versatile editing model PowerPaint [29]. To ensure fairness in qualitative comparison, we fine-tuned the base models of these methods and then applied their respective strategies or operations to generate smoke images. As shown in Figure 8, our method demonstrates superior performance compared to previous approaches. The generated smoke exhibits more reasonable results in terms of color, visibility, and semi-transparency, without issues such as excessive smoke density, distortion, or generation failure. Moreover, our method showcases strong capabilities in preserving the semi-transparent properties of smoke edges and maintaining background integrity, enabling seamless interaction with the background. In terms of generation quality, our method can produce images with a resolution of $1024\times 1024$, providing the smoke detection model with richer details of both the target and the background.
(1) Metrics for image inpainting tasks. Unlike some previous works, we did not use FID (Fréchet Inception Distance) as an evaluation metric for image quality.Unlike many unconditional generation tasks, our problem is an editing-based inpainting setting, where only the masked smoke regions are synthesized while the majority of the image remains unchanged. Therefore, the FID (Fréchet Inception Distance), which relies on global feature statistics, becomes dominated by the preserved background and cannot reliably reflect the quality of the generated smoke. Moreover, FID emphasizes overall distributional diversity, which is less relevant in our application, as the goal of forest fire smoke synthesis is to produce plausible and contextually consistent results rather than maximizing visual diversity. Instead, we adopted commonly used metrics for image generation tasks as our evaluation criteria, including: Peak Signal-to-Noise Ratio (PSNR) [39], Learned Perceptual Image Patch Similarity (LPIPS) [40], Mean Squared Error (MSE), and Structural SIMilarity (SSIM) [41]. Additionally, we employed CLIP Similarity (CLIP Sim) [26] to evaluate the similarity between the generated smoke and the text, thereby assessing the realism of the smoke and the quality of the generated images.
@@
TABLE I: Comparison of generative models. Our method achieves state-of-the-art or competitive results across various metrics for smoke image generation.
@@
(2) Metrics for object detection tasks. We used four metrics to evaluate the performance of models trained on both real and mixed datasets: Precision (Prec), Recall, mAP50, and mAP50-95. Precision measures the proportion of correctly predicted positive instances among all predicted positive instances. Recall measures the proportion of actual positive instances that are correctly identified. mAP50 is used to evaluate the accuracy of object localization by the model at different confidence levels, considering only predictions with an Intersection over Union (IoU) greater than or equal to 0.5. mAP50-95 is an extended version of mAP50, which not only calculates the Average Precision (AP) at an IoU threshold of 0.5 but also computes the AP at multiple IoU thresholds ranging from 0.5 to 0.95 in increments of 0.05, ultimately averaging these AP values to obtain mAP50-95. These metrics help us assess the improvement in the detection model’s performance due to the synthetic data.
TABLE II: Comparison of different detection models. By combining real data with synthetic data, we significantly improved the performance of the smoke detection model.
@@
Yolov12 0.638 0.301 0.765 0.601 0.797 0.492 0.849 0.722
Table I presents the comparison results between our method and previous works, demonstrating the effectiveness of MGFDiffusion in terms of image quality and text alignment. The modified Stable Diffusion, which employs a strategy-based approach, exhibits the poorest image quality. This is because the smoke in the images is generated entirely based on the mask, without considering the interaction between the smoke edges and the background, resulting in either failed smoke generation or overly dense and distorted smoke. PowerPaint performs poorly in generating semi-transparent objects, which we attribute to its training being overly biased toward generating objects with fixed shapes. Simply fine-tuning SD-Inpainting is insufficient to fully utilize the information from the mask and masked image, making it difficult to generate reasonably visible smoke within the mask and causing interference from the model’s prior knowledge. Blended Latent Diffusion shows unstable generation performance, failing to produce smoke in certain scenarios.
TABLE III: Comparison of synthetic data and real data. With an equivalent amount of data, the inclusion of our synthetic data results in more substantial performance gains for the smoke detection model compared to the addition of real data.
@@
Yolov6 0.637 0.330 0.641 0.642 0.759 0.464 0.781 0.677
Fast R-CNN 0.683 0.304 0.602 0.755 0.729 0.383 0.725 0.733
Yolov8 0.675 0.350 0.710 0.665 0.795 0.495 0.790 0.728
RetinaNet 0.710 0.323 0.723 0.707 0.806 0.466 0.829 0.764
@@
RT-DETR1 0.702 0.337 0.758 0.690 0.823 0.514 0.863 0.786
Yolov10 0.670 0.341 0.737 0.636 0.777 0.483 0.815 0.688
RT-DETR2 0.742 0.370 0.752 0.745 0.833 0.523 0.837 0.816
Yolov11 0.680 0.349 0.647 0.725 0.793 0.495 0.811 0.721
Yolov12 0.679 0.361 0.781 0.642 0.797 0.492 0.849 0.722
Figure 9: Effect of modules in network architecture. Using only the mask can lead to a decrease in the ability to interact with the background, while using only the masked image or not employing the Joint Cross-Attention (JCA) mechanism can result in varying degrees of degradation in the generation capability.
In Table II, we demonstrated the effectiveness of our synthetic dataset. We conducted experiments using widely adopted detection models, including Yolov6, Yolov8, Yolov9, Yolov10, and Yolov11. To evaluate the performance of the trained detection models in real-world smoke detection scenarios, we used only real data as the test set. The experimental results indicate that mixing the synthetic dataset with the real dataset leads to significant improvements across all metrics for each detection model. This confirms that our high-quality synthetic dataset can effectively enhance the performance of smoke detection models in real-world environments.
We also compared the effects of replacing the synthetic data with real data during model training, as shown in Table III. It can be observed that, with the same amount of data, incorporating synthetic data leads to greater improvements in the smoke detection model compared to adding real data. We attributed this to the fact that the synthetic smoke data is inherently very close to real data, and through the data filtering process, only high-quality synthetic data is retained. Therefore, adding this portion of high-quality synthetic data to the real dataset yields better results than simply adding more real data.
TABLE IV: Effect of filtering module. The filtering module significantly enhances data quality, and utilizing this high-quality data for training leads to a smoke detection model with superior performance.
@@
Yolov12 0.756 0.438 0.832 0.675 0.797 0.492 0.849 0.722
V-C Ablation Study
Figure 10: Effect of feature extractor. ResNet50 serves as a more stable feature extractor and yields better generation outcomes.
Figure 11: Effect of mask random difference loss. Varying the weights can significantly influence the model’s capability to effectively interact with the background.
We also designed experiments to validate the effectiveness of the multimodal large language model in data filtering. For this purpose, we created two datasets: one using filtered synthetic smoke data and the other using randomly selected synthetic smoke data, both mixed with real data in a 1:1 ratio, with a positive-to-negative sample ratio of 1:1. We trained multiple detection models on these two datasets separately. As shown in Table IV, significant improvements were observed across all four metrics used in our evaluation for different detection models.
VI Conclusion
This paper proposes a comprehensive pipeline for forest fire smoke image generation and detection, which includes the acquisition of a pre-trained diffusion model dataset, a feature extraction module, a novel Mask Random Difference loss function, and a data filtering module based on a Multimodal Large Language Model. By leveraging multiple pre-trained models, we generated a training dataset of over 60,000 samples and trained the model with a newly designed projection layer and cross-attention module. Experiments demonstrate that the proposed MFGDiffusion architecture can generate realistic smoke images, outperforming existing methods in visual aspects such as smoke texture, morphology, and transparency, while also achieving better performance on metrics including PSNR, SSIM, LPIPS, MSE, and ClipSim.
Furthermore, based on the characteristics of the generated smoke, we fine-tuned a Multimodal Large Language Model to filter high-quality datasets. By combining synthetic data with real data, we effectively alleviated the issues of limited availability and insufficient diversity in smoke datasets, significantly improving the performance of smoke detection models.
Although the method effectively improves the realism of smoke generation and the performance of detection models, it has not been extended to the video modality and cannot be used for video-based detection models. Our future work will focus on research into generating natural smoke videos, with particular attention to the fidelity and physical rationality during the smoke generation process.